OpenStudy (anonymous):

Solve the equation? x^2-7x-6=0?

OpenStudy (anonymous):

Your answer is \[\frac{ 7+\sqrt{73} }{ 2 }\] and \[\frac{ 7-\sqrt{73} }{ 2 }\] Use this website: https://mathway.com/ It works well for simple things like this.

OpenStudy (anonymous):

sweet thanx

OpenStudy (anonymous):

No problem!